Suppose the production function of a firm is given ×=0.5lthe power of 1/2*k the power of 1/2 price of labor and capital are given $5 and$10 respectively,and the firm has constant cost out lay of$600





1; find profit maximize level of L and K to employee





2; find MRTS of L to K at optimum

1
Expert's answer
2022-03-15T12:07:31-0400

A)

"x=0.5l^{\frac{1}{2}}k^{\frac{1}{2}}"

Given that Pl=$5 and Pk=$10 and C=$600

"X=0.5l^{\frac{1}{2}}k^{\frac{1}{2}}+\lambda(60-5l-10k)"

"X_l=0.25l^{-\frac{1}{2}}k^{\frac{1}{2}}-5\lambda=0" .......1)

"X_k=0.25l^{\frac{1}{2}}k^{-\frac{1}{2}}-10\lambda=0" .....2)

"X_{\lambda}=60-5l-10k=0" .......3)

"\frac{0.25l^{-\frac{1}{2}}k^{\frac{1}{2}}}{0.25l^{\frac{1}{2}}k^{-\frac{1}{2}}}=\frac{5}{10}"

"\frac{k}{l}=\frac{5}{10}"

"l=2k"

Substituting l in (3)

"600-5(2k)-10k=0"

"600-20k=0"

"\bar{k}=30"

"\bar{l}=2(30)=60"


B)"MRTS_{lk}=\frac{-k}{l}"

At the optimum,

"MRTS_{lk}=\frac{-30}{60}=\frac{-1}{2}"